OpenStudy (anonymous):

The expression -2a(b - 4) is equal to which of the following expressions? -2ab + 8a -2ab + 8 -2ab - 8a 2ab + 8 2ab + 8a

OpenStudy (anonymous):

You have to multiply -2a by each Number/variable in the ( ) parenthesis

OpenStudy (anonymous):

So would my answer be -2a + 8?

OpenStudy (anonymous):

Close...-2a X b = ?

OpenStudy (anonymous):

and -2a x 4 = ?

OpenStudy (anonymous):

I'm confused on how to do -2a x b = -2a x 4 = 8a?

OpenStudy (anonymous):

-2a x b = -2ab -2a x -4 = 8a Numbers you multiply together letters you just increase the power of or if different letters just add to the line Example 2a (4a + 6b -3) = 8a^2 + 12ab -6a 2a x 4a = 8a^2 2a x 6b = 12ab 2a x -3 = -6a